The largest swimming pool in the world is in Casablanca. With a perimeter of 1110 meters this rectangular shaped pool is 30 meters longer than 6 times its width. How many meters wide is it
Answer by sp75(12)

1: Draw a diagram (a simple rectangle will do)
2: Identify what you want to know and give it a name (we want to find width, so let's make its name w)
3: Ask yourself, "What information do I have?"
---> You know that the Perimeter is 1110. Mathematically, you also know that the perimeter is just the sum of all the sides of a rectangle or square. Therefore, the equation for the perimeter (we will call it P) is P = L + L + w + w
(****L is equal to the length of the pool****)
So, if P = 1110 and P = L + L + w + w then 1110 = L + L + w + w
Remember that because we will need it for later.
---> You also know that the pool is 30 meters longer than 6*w
"Longer" is a keyword telling you that the you will need to add 30 to 6*x.
This whole phrase is just another way of saying that the length of the pool is 30 meters longer than six times the width.
So, now we know that L = 30 + (6*w).
Note: Cool. Now we know L, we know P and we have an equation.
Remember P = L + L + w + w?
We have one equation and one unknown (which, in this case, is a recipe for success).
4: Let's solve for w:
1110 = (30 + (6*w) ) + (30 + (6*w)) + w + w
Well, since we don't have any numbers outside of the parenthesis to distribute and the numbers inside of the parenthesis are in their simplest for we can re- write this as
1110 = 30 + 6w +30 + 6w + w + w
Combine terms: 1110 = 60 + 14w
Subtract 60 from both sides: 1050 = 14w
Divide both sides by 14: 75 meters = w